# DreamGaussian: Generative Gaussian Splatting for Efficient 3D Content Creation

### v1

@@ 4.2 Qualitative Comparisons
We first provide qualitative comparisons on image-to-3D in Figure 4. We primarily compare with three baselines from both optimization-based methods (Liu et al., 2023b) and inference-only methods (Liu et al., 2023a; Jun & Nichol, 2023). For all compared methods, we export the generated models as polygonal meshes with vertex color or texture images, and render them under ambient lighting. In terms of generation speed, our approach exhibits a noteworthy acceleration when contrasted with optimization-based methods. Regarding the quality of generated models, our method outperforms inference-only methods especially with respect to the fidelity of 3D geometry and visual appearance. In general, our method achieves a better balance between generation quality and speed, reaching comparable quality as optimization-based methods while only marginally slower than inference-only methods. In Figure 5, we compare the results on text-to-3D. Consistent with our findings in image-to-3D tasks, our method achieves better quality than inference-based methods and faster speed than other optimization-based methods. Furthermore, we highlight the quality of our exported meshes in Figure 6. These meshes exhibit uniform triangulation, smooth surface normals, and clear texture images, rendering them well-suited for seamless integration into downstream applications. For instance, leveraging software such as Blender (Community, 2018), we can readily employ these meshes for rigging and animation purposes.
@@ 4.3 Quantitative Comparisons
Ref. View Consistency $\uparrow$ 3.43 2.33 1.81 4.25
Overall Model Quality $\uparrow$ 3.02 1.86 1.52 3.84
@@ 4.4 Ablation Study
We carry out ablation studies on the design of our methods in Figure 7. We are mainly interested in the generative Gaussian splatting training, given that mesh fine-tuning has been well explored in previous methods (Tang et al., 2023a; Lin et al., 2023). Specifically, we perform ablation on three aspects of our method: 1) Periodical densification of 3D Gaussians. 2) Linear annealing of timestep $t$ for SDS loss. 3) Effect of the reference view loss $\mathcal{L}_{\text{Ref}}$. Our findings reveal that the omission of any of these design elements results in a degradation of the quality of the models generated in the first stage. Specifically, the final Gaussians exhibit increased blurriness and inaccuracies, which further affects the second fine-tuning stage.
@@ 5 Limitations and Conclusion
In this work, we present DreamGausssion, a 3D content generation framework that significantly improves the efficiency of 3D content creation. The key features of our work are: 1) We design the generative Gaussian splatting pipeline which is highly efficient for 3D generation. 2) We propose a mesh extraction algorithm that effectively derives textured meshes from 3D Gaussians. 3) With our texture fine-tuning stage, our work can produce ready-to-use 3D assets with high-quality polygonal meshes from either a single image or text description within a few minutes.
Limitations. We share common problems with previous text-to-3D works: Multi-face Janus problem and baked lighting. Fortunately, it’s possible to address these problems with recent advances in multi-view 2D diffusion models (Shi et al., 2023; Liu et al., 2023c; Zhao et al., 2023) and latent BRDF auto-encoder (Xu et al., 2023b). Besides, the back-view texture generated in our image-to-3D results may look blurry. This can be alleviated with longer stage 2 training.
@@ Acknowledgments
This work is supported by the Sichuan Science and Technology Program (2023YFSY0008), National Natural Science Foundation of China (61632003, 61375022, 61403005), Grant SCITLAB-20017 of Intelligent Terminal Key Laboratory of SiChuan Province, Beijing Advanced Innovation Center for Intelligent Robots and Systems (2018IRS11), and PEK-SenseTime Joint Laboratory of Machine Vision. This project is also funded by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ Appendix A Appendix
A.1 More Implementation Details
@@
Evaluation Settings. We adopt the CLIP-similarity metric (Melas-Kyriazi et al., 2023; Qian et al., 2023; Liu et al., 2023a) to evaluate the image-to-3D quality. A dataset of 30 images collected from previous works (Melas-Kyriazi et al., 2023; Liu et al., 2023a; Tang et al., 2023b; Liu et al., 2023c) and Internet covering various objects is used. We then render $8$ views with uniformly sampled azimuth angles $[0,45,90,135,180,225,270,315]$ and zero elevation angle. These rendered images are used to calculate the CLIP similarities with the reference view, and we average the different views for the final metric. We use the laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^1] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k checkpoint to calculate CLIP similarity. For the user study, we render 360 degree rotating videos of 3D models generated from a collection of 15 images. There are in total 60 videos for 4 methods (Zero-1-to-3 (Liu et al., 2023b), One-2-3-45 Liu et al. (2023a), Shap-E Jun & Nichol (2023), and our method) to evaluate. Each volunteer is shown 15 samples containing the input image and a rendered video from a random method, and ask them to rate in two aspects: reference view consistency and overall model quality. We collect results from 40 volunteers and get 600 valid scores in total.
A.2 More Results
@@
In Figure 8, we show more visualization results of our method. Specially, we compare the mesh output before and after our texture fine-tuning stage. We also compare against a SDS-based mesh fine-tuning method for Zero-1-to-3 (Liu et al., 2023b) noted as $*{}^{*}$ (Tang, 2022). Both stages of our method are faster than previous two-stage image-to-3D methods, while still reaching comparable generation quality.
Our method also support images with non-zero elevations. As illustrated in Figure 10, our method can perform image-to-3D correctly with an extra estimated elevation angle as input. We make sure the random elevation sampling covers the input elevation and at least $[-30,30]$ degree.
@@
In Figure 11, we demostrate the text-to-image-to-3D pipeline (Liu et al., 2023a; Qian et al., 2023). We first apply text-to-image diffusion models (Rombach et al., 2022) to synthesize an image given a text prompt, then perform image-to-3D using our model. This usually gives better results compared to directly performing text-to-3D pipeline, and takes less time to generate. We show more animation results from our exported meshes in Figure 9.
@@
We also illustrate the limitations of our method in Figure 12. Our image-to-3D pipeline may produce blurry back-view image and cannot generate fine details, which looks unmatched to the front reference view. With longer training of stage 2, the blurry problem of back view can be alleviated. For text-to-3D, we share common problems with previous methods, including the multi-face Janus problem and baked lighting in texture images.

### v2

@@ 4.2 Qualitative Comparisons
We first provide qualitative comparisons on image-to-3D in Figure 4. We primarily compare with three baselines from both optimization-based methods (Liu et al., 2023b) and inference-only methods (Liu et al., 2023a; Jun & Nichol, 2023). For all compared methods, we export the generated models as polygonal meshes with vertex color or texture images, and render them under ambient lighting. In terms of generation speed, our approach exhibits a noteworthy acceleration compared to other optimization-based methods. Regarding the quality of generated models, our method outperforms inference-only methods especially with respect to the fidelity of 3D geometry and visual appearance. In general, our method achieves a better balance between generation quality and speed, reaching comparable quality as optimization-based methods while only marginally slower than inference-only methods. In Figure 5, we compare the results on text-to-3D. Consistent with our findings in image-to-3D tasks, our method achieves better quality than inference-based methods and faster speed than other optimization-based methods. Furthermore, we highlight the quality of our exported meshes in Figure 6. These meshes exhibit uniform triangulation, smooth surface normals, and clear texture images, rendering them well-suited for seamless integration into downstream applications. For instance, leveraging software such as Blender (Community, 2018), we can readily employ these meshes for rigging and animation purposes.
@@ 4.3 Quantitative Comparisons
Ref. View Consistency $\uparrow$ 3.48 2.34 1.80 4.31
Overall Model Quality $\uparrow$ 3.11 1.91 1.57 3.92
@@ 4.4 Ablation Study
We carry out ablation studies on the design of our methods in Figure 7. We are mainly interested in the generative Gaussian splatting training, given that mesh fine-tuning has been well explored in previous methods (Tang et al., 2023a; Lin et al., 2023). Specifically, we perform ablation on three aspects of our method: 1) Periodical densification of 3D Gaussians. 2) Linear annealing of timestep $t$ for SDS loss. 3) Effect of the reference view loss $\mathcal{L}_{\text{Ref}}$. Our findings reveal that omission of any design elements results in a degradation of the generated model quality. Specifically, the final Gaussians exhibit increased blurriness and inaccuracies, which further affects the second fine-tuning stage.
@@ 5 Limitations and Conclusion
In this work, we present DreamGausssion, a 3D content generation framework that significantly improves the efficiency of 3D content creation. We design an efficient generative Gaussian splatting pipeline, and propose a mesh extraction algorithm from Gaussians. With our texture fine-tuning stage, we can produce ready-to-use 3D assets with high-quality polygonal meshes from either a single image or text description within a few minutes.
Limitations. We share common problems with previous works: Multi-face Janus problem, over-saturated texture, and baked lighting. It’s promising to address these problems with recent advances in score debiasing (Armandpour et al., 2023; Hong et al., 2023), camera-conditioned 2D diffusion models (Shi et al., 2023; Liu et al., 2023c; Zhao et al., 2023; Li et al., 2023b), and BRDF auto-encoder (Xu et al., 2023b). Besides, the back-view texture generated in our image-to-3D results may look blurry, which can be alleviated with longer stage 2 training.
Ethics Statement
We share common ethical concerns to other 3D generative models. Our optimization-based 2D lifting approach relies on 2D diffusion prior models (Liu et al., 2023b; Rombach et al., 2022), which may introduce unintended biases due to training data. Additionally, our method enhances the automation of 3D asset creation, potentially impacting 3D creative professionals, yet it also enhances workflow efficiency and widens access to 3D creative work.
@@ Acknowledgments
This work is supported by the Sichuan Science and Technology Program (2023YFSY0008), National Natural Science Foundation of China (61632003, 61375022, 61403005), Grant SCITLAB-20017 of Intelligent Terminal Key Laboratory of SiChuan Province, Beijing Advanced Innovation Center for Intelligent Robots and Systems (2018IRS11), and PEK-SenseTime Joint Laboratory of Machine Vision. This study is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ Appendix A Appendix
A.1 Preliminary
Score Distillation Sampling (SDS). SDS was initially introduced by Dreamfusion (Poole et al., 2022), providing a framework that leverages pretrained 2D diffusion models as priors to optimize a parametric image generator. A representative example involves employing a differentiable 3D representation, such as NeRF (Mildenhall et al., 2020), as the image generator:
$\mathbf{x}=g_{\Theta}(p)$ (7)
where $\mathbf{x}$ represents the rendered 2D image from the camera pose $p$, and $g_{\Theta}(\cdot)$ denotes the differentiable rendering function with optimizable NeRF parameters $\Theta$. The SDS formulation is expressed as:
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[w(t)(\epsilon_{\phi}(\mathbf{x};t,e)-\epsilon)\frac{\partial\mathbf{x}}{\partial{\Theta}}\right]$ (8)
where $t\sim\mathcal{U}(0.02,0.98)$ is a randomly sampled timestep, $p$ is a randomly sampled camera pose orbiting the object center, $\mathbf{\epsilon}\sim\mathcal{N}(0,1)$ is a random Gaussian noise, $w(t)=\sigma_{t}^{2}$ is a weighting function from DDPM (Ho et al., 2020), $\epsilon_{\phi}(\cdot)$ is the noise predicting function with a pretrained parameters $\phi$, and $e$ is the text embedding. By optimizing this objective, the denoising gradient $(\epsilon_{\phi}(\mathbf{x};t,e)-\epsilon)$ that contains the guidance information is back-propagated to the rendered image $\mathbf{x}$, which will be further back-propagated to the underlying NeRF parameters $\Theta$ through differentiable rendering (Mildenhall et al., 2020). Therefore, the NeRF can be optimized to form a 3D shape corresponding to the text description.
UV Mapping. UV Mapping is used to project a 2D texture image onto the surface of a 3D polygonal mesh. This requires to map each mesh vertex to a position on the image plane, which is stored as the UV coordinates for each vertex. UV unwrapping (Young, 2021) is employed to automatically compute these UV coordinates given a mesh. Retrieving the texture value at any surface point on a triangle involves barycentric interpolation to calculate the UV coordinate. We utilize NVdiffrast (Laine et al., 2020) for texture mapping and differentiable rendering, facilitating the optimization of the texture image through rendered images.
A.2 More Implementation Details
@@
Mesh Extraction. After extracting the mesh using Marching Cubes (Lorensen & Cline, 1998), we apply isotropic remeshing and quadric edge collapse decimation (Cignoni et al., 2008) to control the mesh complexity. Specifically, we first remesh the mesh to an average edge length of $0.015$, and then decimate the number of faces to $10^{5}$.
@@
Evaluation Settings. We adopt the CLIP-similarity metric (Melas-Kyriazi et al., 2023; Qian et al., 2023; Liu et al., 2023a) to evaluate the image-to-3D quality. A dataset of 30 images collected from previous works (Melas-Kyriazi et al., 2023; Liu et al., 2023a; Tang et al., 2023b; Liu et al., 2023c) and Internet covering various objects is used. We then render $8$ views with uniformly sampled azimuth angles $[0,45,90,135,180,225,270,315]$ and zero elevation angle. These rendered images are used to calculate the CLIP similarities with the reference view, and we average the different views for the final metric. We use the laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^1] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k checkpoint to calculate CLIP similarity. For the user study, we render 360 degree rotating videos of 3D models generated from a collection of 15 images. There are in total 60 videos for 4 methods (Zero-1-to-3 (Liu et al., 2023b), One-2-3-45 Liu et al. (2023a), Shap-E Jun & Nichol (2023), and our method) to evaluate. Each volunteer is shown 15 samples containing the input image and a rendered video from a random method, and ask them to rate in two aspects: reference view consistency and overall model quality. We collect results from 60 volunteers and get 900 valid scores in total.
A.3 More Results
@@
Image-to-3D. In Figure 10, we show more visualization results of our method. Specially, we compare the mesh output before and after our texture fine-tuning stage. We also compare against a SDS-based mesh fine-tuning method for Zero-1-to-3 (Liu et al., 2023b) noted as Zero-1-to-3 (Tang, 2022). Both stages of our method are faster than previous two-stage image-to-3D methods, while still reaching comparable generation quality. Our method also support images with non-zero elevations. As illustrated in Figure 11, our method can perform image-to-3D correctly with an extra estimated elevation angle as input. We make sure the random elevation sampling covers the input elevation and at least $[-30,30]$ degree.
@@
Text-to-image-to-3D. In Figure 13, we demostrate the text-to-image-to-3D pipeline (Liu et al., 2023a; Qian et al., 2023). We first apply text-to-image diffusion models (Rombach et al., 2022) to synthesize an image given a text prompt, then perform image-to-3D using our model. This usually gives better results compared to directly performing text-to-3D pipeline, and takes less time to generate. We show more animation results from our exported meshes in Figure 12.
Text-to-3D with MVDream. In Figure 8, we show text-to-3D results using the multi-view diffusion model MVDream (Shi et al., 2023) as the guidance. The multi-face Janus problem can be significantly mitigated by incorporating camera information to the 2D guidance model. However, it still suffers from over-saturation and unsmooth geometry. We further perform an ablation study on the linear timestep annealing in Figure 9. With the timestep annealing, we find the model converges to a more reasonable shape with the same amount of trianing iterations.
@@
Limitations. We also illustrate the limitations of our method in Figure 14. Our image-to-3D pipeline may produce blurry back-view image and cannot generate fine details, which looks unmatched to the front reference view. With longer training of stage 2, the blurry problem of back view can be alleviated. For text-to-3D, we share common problems with previous methods, including the multi-face Janus problem and baked lighting in texture images.
Figure 8: Text-to-3D results with MVDream (Shi et al., 2023) as the guidance model.
Figure 9: Ablation on timestep annealing for text-to-3D. We use MVDream (Shi et al., 2023) as the guidance model.